### v1

# Confidence-Aware Multi-Field Model Calibration

Abstract.
Accurately predicting the probabilities of user feedback, such as clicks and conversions, is critical for advertisement ranking and bidding. However, there often exist unwanted mismatches between predicted probabilities and true likelihoods due to the shift of data distributions and intrinsic model biases. Calibration aims to address this issue by post-processing model predictions, and field-aware calibration can adjust model output on different feature field values to satisfy fine-grained advertising demands. Unfortunately, the observed samples corresponding to certain field values can be seriously limited to make confident calibrations, which may yield bias amplification and online disturbance. In this paper, we propose a confidence-aware multi-field calibration method, which adaptively adjusts the calibration intensity based on the confidence levels derived from sample statistics. It also utilizes multiple feature fields for joint model calibration with awareness of their importance to mitigate the data sparsity effect of a single field. Extensive offline and online experiments show the superiority of our method in boosting advertising performance and reducing prediction deviations.
Model Calibration, Confidence-aware, Multi-field, Online Advertising
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Computational advertisingccs: Information systems Recommender systems
1. Introduction
Estimating the real probabilities of user actions, such as click-through rate (CTR) and conversion rate (CVR), is a fundamental task in online advertising (Richardson et al., 2007; Chen et al., 2016; Chapelle et al., 2014; Tang et al., 2020). In an ideal system, model-predicted probabilities are expected to reflect the true likelihoods of click and conversion, e.g., the real CTR should match the predicted CTR under sufficiently many ad impressions (Niculescu-Mizil and Caruana, 2005; Lee et al., 2012). Unfortunately, over- or under-estimation phenomena of model predictions are prevalent in real-world recommender systems for various reasons, including the limitation of binary targets (Deng et al., 2021; Bella et al., 2010), model inductive biases (Wenger et al., 2020; Bai et al., 2021; Gawlikowski et al., 2023), and rapid shifts of online data distributions (Ovadia et al., 2019; Huang et al., 2021; Shen et al., 2021; Wang et al., 2022). This discrepancy often causes wasted advertising expenditure on ineffective ads, which damages the ROI of advertising campaigns and user experience (Guo et al., 2022).
Figure 1. An example of multi-field model calibration.
Model calibration post-processes the predicted scores based on the observation of real data to align the gaps between predicted and real probabilities. Extensive studies have been conducted on global calibration, such as binning (Zadrozny and Elkan, 2001; Naeini et al., 2015; Zadrozny and Elkan, 2002) and scaling (Guo et al., 2017; Platt et al., 1999; Kull et al., 2019), which adjust the predicted probabilities according to the posterior information of all observed samples. However, global calibration suffers from the bias counteraction problem, e.g., the overall bias can be very small if scores for one advertiser are over-estimated while those for another advertiser are under-estimated (Pan et al., 2020). Therefore, field-aware calibration is more practical in real-world scenarios, which can calibrate samples with the same value in a certain target field (e.g., ad ID) and alleviate biases in different sample subsets (Wei et al., 2022).
Pioneering studies like NeuCalib (Pan et al., 2020) and AdaCalib (Wei et al., 2022) have focused on field-aware calibration by learning additional calibration neural networks to adjust original model responses. However, the samples corresponding to certain feature values can often be too sparse to learn an unbiased and confident calibrating network. As shown in Fig. 1, the calculated CTR is 0.1 if an ad slot accumulates 50 clicks out of 500 impressions, while the 95% confidence interval of CTR is [0.076, 0.129] (calculated by the Wilson interval formula (Wilson, 1927)). Such a wide range indicates strong uncertainty in field-aware calibration. Furthermore, according to our industrial practice, the volumes of click and impression data w.r.t. most field values are smaller than the example above, which poses a severe challenge to the confidence of calibration models. In addition, different fields may have diverse probability score distributions (Fig. 1 left), and ad platforms may select multiple fields as the business targets. Most existing methods mainly focus on single field calibration, and it is impractical to directly merge multiple fields due to the ultra sparsity of samples satisfying multiple field value conditions.
In this paper, we propose a confidence-aware multi-field calibration method named ConfCalib, which can perform fine-grained calibration on multiple feature fields meanwhile keeping awareness of the confidence degree of field data observation to adjust the calibration intensity adaptively. Compared with methods based on neural networks, ConfCalib, as a non-parametric method, performs more robustly in scenarios with severe data sparsity and has low costs of computation and deployment. Concretely, we regard user feedback on each field value as a binomial distribution and use Wilson intervals to compute confidence-aware deviation degrees of model predictions, where larger deviations indicate greater discrepancies between observed and predicted probabilities. We devise a mapping function to derive a smaller deviation and transform the predicted scores to achieve this new deviation. To further eliminate the influence of data sparsity, we develop a simple yet effective multi-field fusion method to synthesize the calibration results on different fields to enhance calibration performance and better accommodate advertising demands. Extensive offline experiments in different scenarios and the online A/B test on our advertising platform demonstrate the effectiveness of our approach.
The main contributions of this paper are summarized as follows:
  • We propose a confidence-aware model calibration method to perform adaptive calibration on recommendation model predictions based on field-wise data sparsity.
  • We introduce a simple yet effective joint calibration method on multiple fields to improve both calibration performance and robustness to data sparsity.
  • Experiments on different offline datasets and the A/B testing on our advertising platform verify the superiority of our method.
Figure 2. The overall framework of our ConfCalib method.
2. Related Work
Model calibration is a classic problem in machine learning, which is proposed to eliminate the discrepancy between the predicted probabilities of classification models and the true likelihood (Wald et al., 2021; Pleiss et al., 2017; Nixon et al., 2019). It has been widely explored due to the requirement of accurate probability prediction in various applications, such as ad auctions (He et al., 2014; McMahan et al., 2013), weather forecasting (Gneiting and Raftery, 2005; DeGroot and Fienberg, 1983), and personalized medicine (Jiang et al., 2012).
Calibration methods calibrate trained prediction models during training or inference phase. Methods integrated into training usually conduct calibration through well-designed loss functions, including AvUC (Krishnan and Tickoo, 2020), MMCE (Kumar et al., 2018), Focal loss (Mukhoti et al., 2020; Lin et al., 2017) and SB-ECE (Karandikar et al., 2021), etc. Besides, label smoothing, which can be interpreted as a modified primary loss, also serves to model calibration (Müller et al., 2019; Pereyra et al., 2017; Liu et al., 2022). For model calibration in recommendation systems, post-hoc calibration methods are more prevalent. These methods transform the output of trained prediction models in the post-processing phase. According to the characteristics of different mapping functions, existing calibration methods can be divided into three categories: scaling-based, binning-based, and hybrid.
Scaling-based methods (Guo et al., 2017; Mozafari et al., 2018; Kull et al., 2019; Platt et al., 1999; Kweon et al., 2022; Kull et al., 2017; Rahimi et al., 2020; Gupta et al., 2021) fit parametric mapping functions for the predicted scores. Platt Scaling (Platt et al., 1999) adopts a linear mapping, assuming that the scores of each class follow normal distributions with the same variance. Beta calibration (Kull et al., 2017), Gaussian calibration (Kweon et al., 2022) and Gamma calibration (Kweon et al., 2022) further assume more complicated data distributions and derive corresponding score mapping functions to adapt more general situations. Dirichlet calibration (Kull et al., 2019) generalizes Beta calibration to the multi-class classification by assuming a Dirichlet distribution. The above methods are limited to a few occasions due to the specific distribution assumptions. Temperature Scaling (Guo et al., 2017) considers that the miscalibration of models comes from the overfitting of negative log-likelihood (NLL) during training, then proposes a temperature multiplier for logits of all classes to reduce test NLL and eliminate miscalibration. Although these methods have solid theoretical validity, they are difficult to deal with complex practical scenarios.
Binning-based methods (Zadrozny and Elkan, 2002; Menon et al., 2012; Neelon and Dunson, 2004; Zadrozny and Elkan, 2001; Naeini et al., 2015) divide samples into several bins after sorting them by predicted scores. Histogram Binning (Zadrozny and Elkan, 2001) adopts an equal-frequency binning and directly uses the mean positive rate in each bin as the calibrated scores. Bayesian Binning (Naeini et al., 2015) considers multiple binning schemes and combines them using a derived Bayesian score to yield more robust calibrated predictions. To maintain the original order of model predictions, Isotonic Regression (Zadrozny and Elkan, 2002) adjusts the binning bounds by minimizing the residual between calibrated predictions and labels, hence ensuring monotonicity of the posterior probability in bins.
Hybrid methods (Kumar et al., 2019; Zhang et al., 2020; Deng et al., 2021; Huang et al., 2022; Pan et al., 2020; Wei et al., 2022) combine scaling and binning to fit various mapping functions for different sample subsets, which allows a more precise calibration than single scaling or binning methods. Scaling-binning (Kumar et al., 2019) first fits a mapping function for the sorted samples and then employs histogram binning calibration on the mapped scores. Smoothed Isotonic Regression (SIR) (Deng et al., 2021) uses isotonic regression on bins and fits a monotonic piece-wise linear calibration function. Multiple Boosting Calibration Trees (MBCT) (Huang et al., 2022) optimizes the binning scheme through a tree-like division along feature fields and adopts linear scaling on each tree node individually. Recently, some works have focused on field-wise calibration, which aims to reduce the calibration error in a specific feature field. Neural Calibration (Pan et al., 2020) learns a neural isotonic line-plot scaling function and an auxiliary network directly taking features as input to capture feature field information. AdaCalib (Wei et al., 2022) extends the single global mapping function of Neural Calibration to different ones for each value of a certain feature field and dynamically adjusts the number of bins by neural networks. These neural methods may suffer from data sparsity in specific application scenarios because networks are difficult to train. Additionally, existing field-wise calibration methods simply consider one field and ignore potential relations among fields, meaning there is still a lack of research on multi-field calibration methods.
3. Methodology
The overall framework of our multi-field confidence-aware calibration method is shown in Fig. 2. It first performs confidence-aware calibration on multiple target feature fields independently and then fuses the scores calibrated on different feature fields into a unified one. The details of our approach are introduced as follows.
3.1. Problem Definition
First, we briefly introduce the problem setting in this paper, namely field-level calibration on the binary classification problem. Suppose we have a trained binary classification model $f(\cdot)$ to provide predictions $\hat{p}=f(x)$ on a dataset $\{x_{i},y_{i}\}_{i=0}^{N}$, where $x_{i}$ denotes the features with multiple fields and $y_{i}$ denotes the binary label. We generally consider biases on the level of sample subsets since the true likelihood of a specific sample is unobservable. For a specific sample subset, the average predicted probability usually deviates from the observed positive sample rate, thus requiring calibrations. Assume there are $K$ target feature fields and the value of each field is denoted as $z_{j}~{}(j=1,2,\dots,K)$, the calibration model fits a transformation function $g(\cdot;z_{1},\dots,z_{K})$ and gives calibrated probability $\hat{p}_{\text{calib}}=g(\hat{p};z_{1},\dots,z_{K})$ according to sample statistics on target fields. The goal of field-level calibration on a certain field is to minimize the field-level calibration error, such as Field-RCE (Pan et al., 2020). For multi-field calibration, we can employ metrics like ECE (Naeini et al., 2015) and MVCE (Huang et al., 2022) to measure the global calibration error. The above metrics will be detailed in Section 4.1.4.
3.2. Confidence-aware Calibration
Figure 3. The curves of calibrated scores w.r.t. different predicted scores under different numbers of observed samples.
Next, we introduce our confidence-aware calibration method. In practical scenarios, field-level calibration faces severe data sparsity issues due to the limited time window for data collection and the intrinsic sparsity of certain user feedback, such as conversions. Consequently, calibrations on fields with numerous feature values are neither accurate nor confident. The main problem caused by data sparsity is that the observed probability is not credible enough to reflect the true data distribution. In this case, complete trust in limited observations will likely result in overfitting. Therefore, it is necessary to preserve some trust in model predictions while referencing observations under data sparsity.
Actually, many types of user feedback, like clicks and conversions, can be naturally modeled as binomial distributions. Given the observed positive sample ratio $p$ and the number of total samples $n$, we can then use a confidence interval formula to compute the confidence interval ($p_{-}$, $p_{+}$) of the real feedback probability under a certain confidence level. Conversely, given a certain predicted feedback probability, we can obtain the corresponding confidence level by considering the predicted value as the bound of the confidence interval. In other words, there is a mapping between the predicted probability and the confidence level, which is the core of our confidence-aware calibration method.
The most commonly used confidence interval formula is the normal interval. However, due to the limitation of the central limit theorem, it is unreliable with a small sample size or an extreme probability close to 0 or 1, which is quite common for user feedback prediction in recommender systems. Therefore, we adopt the Wilson confidence interval (Wilson, 1927), which can be safely employed under the special circumstances above. The formula is
(1) $\small(p_{-},p_{+})\approx{\frac{1}{~{}1+{\frac{\,z^{2}\,}{n}}~{}}}\left({p}+{\frac{\,z^{2}\,}{2n}}\right)~{}\pm~{}{\frac{z}{~{}1+{\frac{z^{2}}{n}}~{}}}{\sqrt{{\frac{\,{p}(1-{p})\,}{n}}+{\frac{\,z^{2}\,}{4n^{2}}}~{}}},$
where $z$ is the deviation score (e.g., $z=1.96$ for 95% confidence).
In order to perform field-level calibration, our approach operates at the level of sample subsets, which can be obtained by splitting samples based on different values of the target feature field. Besides, field-wise binning [^1] Binning is optional in our method since it introduces heavier data sparsity., i.e., further dividing each subset into several bins by order of their predicted scores, is optional. For each sample subset, we need the statistics, including the number of samples $n$, the ratio of positive samples $p$, and the average predicted score $\hat{p}$. By substituting these variables into Eq. (1), we can solve the value of the deviation score $z\in[0,+\infty)$ using numerical algorithms, such as the bisection method. A larger $z$ indicates greater discrepancies between model predictions and current observations. To calibrate the predictions, we propose to use a non-linear transformation function $g(z)$ to reduce the deviation. Intuitively, this function should satisfy the following requirements: (1) passing through the origin: $g(0)=0$; (2) monotonic: $\forall z_{1}\leq z_{2},g(z_{1})\leq g(z_{2})$; (3) bounded: $\exists C\in\mathbb{R}^{+},g(+\infty)\leq C$. Thus, we devise a variant of the sigmoid function as follows:
(2) $\small g(z)=\lambda\cdot\left(\frac{2}{1+e^{-z/2}}-1\right),z\in[0,+\infty),$
where $\lambda$ is a hyperparameter used to provide an upper bound and control the calibration intensity manually for large z-scores, i.e., deviations. A large calibration intensity means more belief in observations and less belief in original predictions.
Obviously, we have $g(+\infty)=\lambda$ so that large deviations are calibrated to a relatively small one (e.g., the deviation corresponding to 95% confidence level if $\lambda=1.96$), and minor deviations are reduced almost linearly. By replacing the original $z$ in Eq. (1) with the transformed value $g(z)$, we can obtain the calibrated average probability score $\hat{p}^{\prime}$. In order to ensure that $\hat{p}^{\prime}$ is located between the observed probability $p$ and the predicted probability $\hat{p}$, we use the positive branch if $\hat{p}$ is greater than $p$. Otherwise, we use the negative one. In this way, given a test sample located in the same subset, we compute a field-wise calibrated score $p_{\text{calib}}$ by scaling its original predicted probability $p_{\text{pred}}$, i.e.,
(3) $\small p_{\text{calib}}=\frac{\hat{p}^{\prime}}{\hat{p}}p_{\text{pred}}.$
Fig. 3 shows the curves of calibrated probabilities w.r.t variant predicted probabilities under different observed sample numbers and positive sample rates, taking the conversion rate prediction task as an example. We uniformly set $\lambda=2$ for the z-score transformation function. The curves demonstrate that the predicted probability is calibrated towards the observed one, and the proximity to the observation increases as the number of samples increases.
Overall, our confidence-aware method adjusts model-predicted probabilities by resetting the confidence level, which considers the observation and prediction together to avoid violent value changes and overfitting limited observations. Consequently, it keeps robust when handling severe sample sparsity, such as circumstances with few observed samples and zero positive samples.
3.3. Multi-field Joint Calibration
Previous field-level calibration methods only calibrate and evaluate errors in a certain field, ignoring other fields. Actually, the field-level calibration error in different fields may vary greatly even for the same dataset, indicating a natural discrepancy of prediction biases in different fields. In addition, the biases in different fields, as reflections of the global bias from different perspectives, have a certain degree of commonality. Therefore, the data sparsity problem faced by a certain field is expected to be alleviated through calibrations on other fields. Therefore, simultaneously capturing biases of multiple fields can be significant to eliminate overall prediction bias and improve calibration performance.
A naive way to extend calibration from a single field to multiple fields is to split sample subsets based on the combination of feature values of all the fields. Nevertheless, it is infeasible for the severer data sparsity since fewer samples satisfy multiple field value conditions. This means prior field-wise calibration methods for a single field cannot be directly extended to multi-field calibration.
On the contrary, our method demonstrates better extensibility among fields. Note that the confidence-aware calibration of a certain field in Section 3.2 is actually a scaling operation for a specific sample and thus can be easily generalized to a multi-field joint calibration method. Assume that there are $K$ target fields to calibrate. Given a sample, we will get $K$ scaling multipliers $m_{1},m_{2},\dots,m_{K}$, i.e. ${\hat{p}^{\prime}}/{\hat{p}}$ of each field. Then, a natural manner to adaptively fuse these multipliers is the weighted geometric mean,
(4) $p_{\text{calib}}=m_{1}^{w_{1}}m_{2}^{w_{2}}\cdots m_{K}^{w_{K}}\cdot p_{\text{pred}},$
where $w_{1},w_{2},\dots,w_{K}$ is weights satisfying $\sum_{i=1}^{K}{w_{i}}=1$. Considering the relatively small computation cost, the fusion weights can be precisely obtained by grid search. This method allows us to conduct joint calibration based on the marginal distribution of each field independently instead of dealing with the joint distribution, thus avoiding further sparsity of samples.
3.4. Online Deployment
Fig. 4 shows the online deployment diagram of ConfCalib in our advertising platform. On the offline side, the prediction model’s training and ConfCalib’s fitting are paralleled and completely independent. At fixed intervals, the calibration module pulls the latest online samples from the log database and refits calibration curves. This interval is usually set relatively shorter (e.g., 30 minutes) than the prediction model update period to better capture the dynamic shift of online data. On the online side, prediction and calibration are conducted in sequence. Firstly, the prediction model makes predictions for candidate objects, then the calibration model calibrates the predicted probabilities and outputs the calibrated probabilities for subsequent use in the platform.
Recent methods designed for advertising platforms, such as NeuCalib and AdaCalib, involve auxiliary neural networks. These calibration models are coupled with original prediction models during the training phase, thus increasing the complexity of online deployment. On the contrary, our ConfCalib relies entirely on posterior probabilities of online samples and fits parameters without training networks. Therefore, it can be directly inserted into the pipeline without interfering with the training stage of prediction models, resulting in lower deployment costs.
Figure 4. Online deployment diagram of ConfCalib.
4. Experiments
4.1. Experimental Setup
4.1.1. Settings
The pipeline of our experiment is as follows. Each dataset is conventionally split into three parts in chronological order, i.e., training, validation, and test set. Firstly, we train a base neural network on the training set as the prediction model and fix it. Then, the base model makes predictions for validation set samples to fit (or train) calibration models. Finally, we use the test set to evaluate the performance of calibration methods.
For all the datasets and methods, the prediction model is a fully connected network with four layers. Hyperparameters of comparative methods, like the number of bins, are set according to the chosen value in the corresponding papers if provided. Otherwise, we tune the hyperparameters and select the optimal values. Considering the strong randomness of neural methods, we conduct five random experiments on each of them and report means and standard deviations of the results. For our method we set the number of bins as {50, 1, 10} and $\lambda$ in the z-score mapping function as {1.0, 0.2, 1.6}, corresponding to the following three datasets.
Table 1. Detailed statistics of the three datasets.
Dataset Task #Training #Validation #Test
Avazu CTR 28.3M 4.04M 8.08M
AliExpress CTCVR 62.2M 10.4M 31.2M
Industrial CVR 27.5M 3.8M 7.5M
Table 2. The performance of different methods in terms of prediction accuracy and miscalibration degrees. $\uparrow$” means higher is better, $\downarrow$” means lower is better.
Dataset Type Method AUCnormal-$\uparrow$ LogLossnormal-$\downarrow$ Field-RCEnormal-$\downarrow$ Multi-field-RCEnormal-$\downarrow$ ECEnormal-$\downarrow$ MVCEnormal-$\downarrow$
Avazu (CTR) No Calib. N/A 0.7545 0.3718 0.2295 0.1837 0.0123 0.0117
Binning HistoBin (Zadrozny and Elkan, 2001) 0.7543 0.3714 0.2034 0.1457 0.0149 0.0038
IsoReg (Zadrozny and Elkan, 2002) 0.7545 0.3713 0.1997 0.1432 0.0082 0.0039
Scaling Naive 0.7545 0.3713 0.1874 0.1381 0.0056 0.0038
PlattScaling (Platt et al., 1999) 0.7545 0.3712 0.1868 0.1383 0.0058 0.0039
GaussCalib (Kweon et al., 2022) 0.7545 0.3712 0.1969 0.1470 0.0058 0.0041
GammaCalib (Kweon et al., 2022) 0.7545 0.3712 0.1984 0.1480 0.0058 0.0041
Hybrid SIR (Deng et al., 2021) 0.7545 0.3717 0.2316 0.1849 0.0122 0.0116
NeuCalib (Pan et al., 2020) 0.7554$\pm$0.0008 0.3709$\pm$0.0006 0.1491$\pm$0.0209 0.1091$\pm$0.0225 0.0110$\pm$0.0023 0.0042$\pm$0.0021
AdaCalib (Wei et al., 2022) 0.7569$\pm$0.0003 0.3696$\pm$0.0002 0.1650$\pm$0.0155 0.1308$\pm$0.0247 0.0099$\pm$0.0012 0.0065$\pm$0.0029
ConfCalib 0.7575 0.3693 0.1435 0.0967 0.0037 0.0015
AliExpress (CTCVR) No Calib. N/A 0.8773 0.0046 0.3274 0.3273 2.085e-4 2.085e-4
Binning HistoBin 0.8752 0.0047 0.1417 0.1411 8.688e-5 8.742e-5
IsoReg 0.8770 0.0046 0.1397 0.1394 9.811e-5 8.568e-5
Scaling Naive 0.8773 0.0046 0.1431 0.1429 8.379e-5 8.450e-5
PlattScaling 0.8773 0.0045 0.1406 0.1403 8.597e-5 8.658e-5
GaussCalib 0.8773 0.0045 0.1433 0.1430 8.784e-5 8.836e-5
GammaCalib 0.8773 0.0045 0.1404 0.1401 8.538e-5 8.600e-5
Hybrid SIR 0.8773 0.0046 0.3299 0.3298 2.102e-4 2.102e-4
NeuCalib 0.8773$\pm$0.0003 0.0046$\pm$5e-7 0.1499$\pm$0.0480 0.1497$\pm$0.0480 1.068e-4$\pm$3.236e-5 8.617e-5$\pm$3.300e-5
AdaCalib 0.8759$\pm$0.0004 0.0046$\pm$5e-8 0.1459$\pm$0.0383 0.1458$\pm$0.0375 1.045e-4$\pm$3.131e-5 8.003e-5$\pm$3.316e-5
ConfCalib 0.8778 0.0045 0.1360 0.1360 8.352e-5 8.425e-5
Industrial (CVR) No Calib. N/A 0.8226 0.0868 0.3657 0.3015 0.0022 0.0008
Binning HistoBin 0.8223 0.0868 0.3838 0.3145 0.0017 0.0015
IsoReg 0.8226 0.0867 0.3779 0.3079 0.0036 0.0015
Scaling Naive 0.8226 0.0868 0.3555 0.2811 0.0020 0.0016
PlattScaling 0.8226 0.0868 0.3929 0.3190 0.0018 0.0015
GaussCalib 0.8226 0.0867 0.3606 0.2916 0.0017 0.0015
GammaCalib 0.8226 0.0867 0.3756 0.3047 0.0018 0.0015
Hybrid SIR 0.8226 0.0868 0.3555 0.2811 0.0020 0.0016
NeuCalib 0.8271$\pm$0.0020 0.0866$\pm$0.0002 0.2856$\pm$0.0366 0.2330$\pm$0.0234 0.0012$\pm$0.0006 0.0007$\pm$0.0004
AdaCalib 0.8192$\pm$0.0005 0.0875$\pm$0.0001 0.3234$\pm$0.0231 0.2307$\pm$0.0295 0.0010$\pm$0.0003 0.0007$\pm$0.0004
ConfCalib 0.8303 0.0860 0.2728 0.2243 0.0007 0.0007
4.1.2. Datasets
We conduct experiments on our industrial dataset, as well as two public datasets, including Avazu  [^2] https://kaggle.com/competitions/avazu-ctr-prediction and AliExpress (Li et al., 2020). The data statistics are shown in Table 1.
Avazu is a popular dataset on mobile ad click-through rate (CTR) prediction with 40.4 million samples in 10 days ordered by the impression time. We split the dataset by day in a ratio of 7:1:2 for training, validation, and test. For our multi-field joint calibration method, the target fields are set to ”site_id” (3480) [^3] Represent the number of possible values of the corresponding field., ”app_id” (4864) and ”app_domain” (305). For other methods considering a single field, the target field is set to ”site_id”.
AliExpress is a dataset for multi-task learning of click-through rate and conversion rate prediction in e-commerce. Here, we choose the data from four countries except Russia and only use the conversion label to build a single click-through conversion rate (CTCVR) prediction task. The training, validation, and test set are split in a ratio of 6:1:3, totaling 103.8 million samples. The timestamp of the test set is strictly later than the other two sets. For our method, the target fields are ”country” (4) and an anonymous field, ”feature_1” (10). For other methods, the target field is ”feature_1”.
For the industrial data, we collect 38.8M clicked samples in 10 consecutive days from the log of our ad platform to build a post-click conversion rate (CVR) prediction task with a split ratio of 7:1:2 by day. For our method, the target fields for calibration are ”advertiser_id” (529), ”app_size” (82), and ”app_id” (56). For other methods, the target field is ”advertiser_id”.
The ratios of validation sets are relatively small while splitting, which is more in line with the short time window for data collection during online running.
4.1.3. Comparative Methods
We choose several representative methods of the three categories of methods for comparison. The scaling-based methods include naive scaling (multiply all predicted scores by a value $k$ to make their mean equal to the positive sample ratio), Platt Scaling (PlattScaling) (Platt et al., 1999), Gaussian Calibration (GaussCalib) (Kweon et al., 2022) and Gamma Calibration (GammaCalib) (Kweon et al., 2022). The binning-based methods include Histogram Binning (HistoBin) (Zadrozny and Elkan, 2001) and Isotonic Regression (IsoReg) (Zadrozny and Elkan, 2002). The hybrid methods include Smoothed Isotonic Regression (SIR) (Deng et al., 2021), Neural Calibration (NeuCalib) (Pan et al., 2020) and AdaCalib (Wei et al., 2022), with the latter two methods based on neural networks. The number of bins for NeuCalib and AdaCalib is set according to the initial settings when proposed.
4.1.4. Evaluation Metrics
Referring to prior methods, we continue to use field-level relative calibration error (Field-RCE) (Pan et al., 2020) to evaluate the calibration error on a certain field, which formula is
[…]
where $\mathcal{D}$ is the dataset, $\mathcal{Z}$ is the possible value set of the target field and $\mathcal{D}_{z}$ is the subset of samples with the target field value $z$ satisfying $\sum_{z\in\mathcal{Z}}{|\mathcal{D}_{z}|}=|\mathcal{D}|$. Intuitively, a lower Field-RCE indicates lower biases along the target field. Furthermore, we use the mean Field-RCE on multiple fields (Multi-field-RCE) to measure the multi-field calibration performance.
Meanwhile, we adopt expected calibration error (ECE) (Naeini et al., 2015) and multi-view calibration error (MVCE) (Huang et al., 2022) to evaluate the global calibration error. ECE arranges all predicted values in order and divides them equally into several subsets, i.e., a binning operation. Then it is calculated as
(6) $\small\text{ECE}=\frac{1}{|\mathcal{D}|}\sum_{t=1}^{T}{{\left|\sum\nolimits_{(x_{i},y_{i})\in\mathcal{D}_{t}}{(y_{i}-\hat{p}_{i})}\right|}},$
where $T$ is the number of bins and $\mathcal{D}_{t}$ is the subset of samples in the $t$-th bin. MVCE iteratively shuffles the dataset and calculates ECE until the accumulative mean ECE converges. The formula is
(7) $\small\text{MVCE}=\frac{1}{RT}\sum_{r=1}^{R}\sum_{t=1}^{T}{\frac{\left|\sum_{(x_{i},y_{i})\in\mathcal{D}_{r,t}}{(y_{i}-\hat{p}_{i})}\right|}{|\mathcal{D}_{r,t}|}},$
where $R$ is the number of shuffle operations, $T$ is the number of bins, and $\mathcal{D}_{r,t}$ is the subset of samples in the $t$-th bin after the $r$-th shuffle. MVCE allows us to inspect the calibration error from a more global perspective beyond the level of feature fields, thereby highlighting the effectiveness of multi-field calibration. The above calibration error metrics focus on different aspects, thus allowing us to examine the performance of each method more comprehensively.
Additionally, we also report AUC and LogLoss to evaluate the ranking performance after calibration since our method and main comparative methods are not globally isotonic.
4.2. Results
Figure 5. Results of ablation studies on several variants of ConfCalib.
4.2.1. Main Results
The main results on each dataset are shown in Table 2. Our method shows better calibration performance compared with the competing methods. It achieves lower calibration errors on both field-wise and global metrics. Besides, our method also achieves higher AUC in all the tasks, indicating improvement in ranking ability in contrast to the uncalibrated model and other methods that may change the original order.
By contrast, the performance of scaling and binning-based comparative methods is unsatisfactory. These methods reduce global errors to some extent through global calibration, but the performance is not outstanding. Meanwhile, they are basically powerless against field-wise calibration errors.
Field-level calibration methods with neural networks such as NeuCalib and AdaCalib achieve relatively low calibration errors on the single target field among comparative methods. However, they are still inferior to our method, especially on multi-field metrics like Multi-field-RCE and global metrics like ECE, demonstrating the effectiveness of our multi-field joint calibration. Moreover, the results of the two neural methods show large variance, which means significantly unstable learning processes on small calibration datasets. On the contrary, our method only relies on posterior statistical information and maintains a stabler calibration effect.
It is noteworthy that results of the neural method AdaCalib on industrial data and AliExpress indicate that it may be unable to learn effectively in some cases, resulting in AUC even lower than the uncalibrated one. This phenomenon is likely caused by severe data sparsity. After sequential divisions of the feature value split and the binning operation, the number of total and positive samples within each subset is very limited. Thus, it is difficult for AdaCalib to learn the network parameters of corresponding parts of the piece-wise calibration curves. Our method works well in the same situation, benefiting from the awareness of confidence and the multi-field fusion approach designed to deal with sparse data.
4.2.2. Ablation Study
We report the results of several variants of our ConfCalib on Avazu CTR task in Fig. 5. Concretely, we respectively remove the confidence-aware calibration with Wilson interval, the nonlinear z-score mapping function (replaced by a single linear mapping), and the multi-field joint calibration from the original ConfCalib. The results show that all variants have varying degrees of increase in calibration errors and decrease in AUC, which can demonstrate the effectiveness of each key point.
The confidence-aware calibration partly preserves confidence in original predictions while calibrating based on observed posterior probabilities, which can avoid overfitting on limited validation samples in case the bias direction in the test dataset is opposite. In other words, it can effectively mitigate the impact of real-time data shifts during online serving. Although we cannot simulate frequent data shifts using a fixed test dataset, there is still an evident decline in AUC and an increase in Field-RCE after removing the confidence-aware calibration.
The elaborate z-score mapping function dynamically adjusts the calibration intensity according to the deviations between predicted and observed probabilities. For samples with large deviations, it introduces an upper bound to force the calibrated result back to a high enough confidence level. After replacing the mapping function with a single linear function $g(z)=\lambda z$, AUC and all calibration error metrics significantly deteriorate. The main reason is that the unbounded linear function chooses insufficient calibration strength for samples with large deviations.
The multi-field joint calibration simultaneously considers the biases across multiple fields and conducts joint calibration with optimized weights of each field. Compared to calibration on a single field, the joint calibration on multiple fields significantly improves both field-wise and global calibration errors, as well as AUC.
Figure 6. Results of robustness analysis of different data sparsity levels.
Table 3. Results of recalibration on comparative methods.
Method AUCnormal-$\uparrow$ Field-RCEnormal-$\downarrow$ Multi-field-RCEnormal-$\downarrow$ ECEnormal-$\downarrow$ MVCEnormal-$\downarrow$
GaussCalib
0.7545
0.1969
0.1470
0.0058
0.0041
+ConfCalib 0.7577 0.1417 0.1105 0.0053
0.0050
GammaCalib
0.7545
0.1984
0.1480
0.0058
0.0041
+ConfCalib 0.7577 0.1430 0.1113 0.0053
0.0050
NeuCalib
0.7554
0.1491
0.1091
0.0110
0.0042
+ConfCalib 0.7558 0.1446 0.0994 0.0099 0.0030
AdaCalib
0.7569
0.1650
0.1308
0.0099
0.0065
+ConfCalib
0.7569
0.1535 0.1069 0.0072 0.0024
4.3. Recalibration on Comparative Methods
Part of the comparative methods are learning-based, such as GaussCalib, GammaCalib, NeuCalib, and AdaCalib. It can be seen from the results in Table 2 that the performances of these methods are not satisfying for remaining high calibration errors or reducing AUC. Since our method does not require extra model training and has a low deployment cost, it can recalibrate the calibrated output of other methods for further improvement.
We experiment on Avazu CTR task. In detail, we first independently fit our method and the comparative methods to be recalibrated on the validation set. Then, the comparative methods make calibrated predictions for the test set. Finally, our method conducts recalibration on the calibrated outputs. Table 3 shows that recalibration using our ConfCalib achieves further improvement compared with the original results of other calibration methods. Profiting from the multi-field joint calibration, the recalibration significantly reduces Multi-field-RCE and the global calibration error MVCE. Besides, there is a large increase in AUC for scaling-based methods, including GaussCalib and GammaCalib, which strictly keep the order of the original predictions.
4.4. Robustness Analysis against Data Sparsity
As described in Section 3.2, one main advantage of ConfCalib is its robustness against data sparsity. Due to the limited time window for online data collection, the number of samples available to fit calibration models can be very few. Under such circumstances, methods based on neural networks can lead to inferior effects, although they may behave well if trained on sufficient samples.
Observing the results on Avazu in Table 2, we can find that the neural methods NeuCalib and AdaCalib both achieve lower calibration errors compared with the original predictions. Here, we reduce the number of validation samples in Avazu with various sample rates to simulate scenarios with varying degrees of data sparsity. The calibration performances w.r.t different sparsity degrees are shown in Fig. 6. As the number of samples for fitting the calibration model decreases, the performances of NeuCalib, AdaCalib, and ConfCalib all deteriorate in general. Fluctuations in the results of the two neural methods are intense, indicating that they are susceptible to data sparsity. However, our ConfCalib is little affected and exhibits strong robustness against data sparsity.
4.5. Analysis of Multi-field Fusion Weights
The multi-field joint calibration of ConfCalib obtains fusion weights for multiple target fields by minimizing the calibration error on single or multiple fields. Here, we select three target fields, ”ad_type”, ”consume_frequency”, and ”price” in our industrial dataset, and optimize Field-RCE of each field and Multi-field-RCE using grid search to acquire four groups of fusion weights separately, as shown in Fig. 7. Each column takes Field-RCE of the corresponding field as the optimization object. We can see that each field makes effects with different weights when optimizing Multi-field-RCE of all three fields. When optimizing Field-RCE on a single field, the weights of other fields are not zero, indicating that joint calibration on multiple fields can also reduce the single-field calibration error. The weight of ”ad_type” is generally the highest, suggesting that this field can better reflect global biases. Nevertheless, when Field-RCE on ”ad_type” is minimized, the weight of the other field ”consume_frequency” is the highest instead. This phenomenon indicates that calibration on a single field may not be sufficient, while calibration on other fields can provide effective assistance.
Figure 7. Fusion weights in multi-field joint calibration.
4.6. Hyperparameter Analysis
Our method has a single hyperparameter $\lambda$ located in the z-score mapping function to control the calibration strength manually. Its main effect is to select a relatively higher confidence level when observed samples are adequate, thus making the calibrated prediction closer to the observation. Here, we discuss the influence of different settings of $\lambda$. In detail, we set $\lambda\in\{0,0.2,0.4\dots,3.0\}$ and perform experiments on the industrial dataset. The results are shown in Fig. 8. Overall, the fluctuation of field-wise calibration errors like Field-RCE and Multi-field-RCE is within a small range as $\lambda$ changes. For global calibration error metrics, the change of ECE is relatively violent when $\lambda$ is large, while MVCE remains almost unchanged throughout. When $\lambda=0$, the awareness of confidence is removed, which means we disregard the model prediction and believe entirely in the observation. We achieve a local optimum at this point. As $\lambda$ increases, the calibration intensity gradually decreases, and the global optimal $\lambda$ is around 1.6, considering all the calibration errors.
Figure 8. Calibration errors w.r.t hyperparameter $\lambda$.
4.7. Online A/B Test
We conduct one week’s online A/B test separately on two business scenarios to evaluate the effectiveness of ConfCalib in practical applications. In each scenario, we randomly split users into two groups, each with tens of millions of users everyday. One group receives recommendations by ConfCalib, while the other one gets recommendations by a well-crafted baseline. On an industrial advertising platform, ConfCalib achieves a relative improvement of 2.42% in CVR compared with the baseline. On the Top-grossing of a popular Android app store, ConfCalib obtains relative improvement of 32.6% and 49.1% in CTR and the revenue over the baseline. The data distribution shifts frequently on the Top Grossing scenario, and ConfCalib can better adapt this variability to achieve a significant improvement.
5. Conclusion
In this paper, we propose a confidence-aware multi-field model calibration method, ConfCalib, to address the common phenomenon of data sparsity in model calibration on advertising platforms. ConfCalib can dynamically adjust the calibration intensity based on the number of observed samples and posterior probabilities while considering the biases along multiple feature fields for joint calibration to capture various bias patterns and further alleviate the impact of data sparsity. We conduct sufficient comparative experiments on public and industrial datasets of different user action prediction tasks, and the results demonstrate that ConfCalib outperforms prior methods and exhibits strong robustness against data sparsity. ConfCalib is currently running on an industrial advertising platform and a popular Android app store, and has achieved continuous online revenue improvement.

### v2

@@
 Accurately predicting the probabilities of user feedback, such as clicks and conversions, is critical for advertisement ranking and bidding.
-However, there often exist unwanted mismatches between predicted probabilities and true likelihoods due to the shift of data distributions and intrinsic model biases.
+However, there often exist unwanted mismatches between predicted probabilities and true likelihoods due to the rapid shift of data distributions and intrinsic model biases.
 Calibration aims to address this issue by post-processing model predictions, and field-aware calibration can adjust model output on different feature field values to satisfy fine-grained advertising demands.
 Unfortunately, the observed samples corresponding to certain field values can be seriously limited to make confident calibrations, which may yield bias amplification and online disturbance.
-In this paper, we propose a confidence-aware multi-field calibration method, which adaptively adjusts the calibration intensity based on the confidence levels derived from sample statistics.
-It also utilizes multiple feature fields for joint model calibration with awareness of their importance to mitigate the data sparsity effect of a single field.
+In this paper, we propose a confidence-aware multi-field calibration method, which adaptively adjusts the calibration intensity based on confidence levels derived from sample statistics.
+It also utilizes multiple fields for joint model calibration according to their importance to mitigate the impact of data sparsity on a single field.
 Extensive offline and online experiments show the superiority of our method in boosting advertising performance and reducing prediction deviations.
@@
 Therefore, field-aware calibration is more practical in real-world scenarios, which can calibrate samples with the same value in a certain target field (e.g., ad ID) and alleviate biases in different sample subsets (Wei et al., 2022).
-Pioneering studies like NeuCalib (Pan et al., 2020) and AdaCalib (Wei et al., 2022) have focused on field-aware calibration by learning additional calibration neural networks to adjust original model responses.
-However, the samples corresponding to certain feature values can often be too sparse to learn an unbiased and confident calibrating network.
+Pioneering studies like NeuCalib (Pan et al., 2020) and AdaCalib (Wei et al., 2022) have focused on field-aware calibration by learning additional calibration neural networks to adjust original model predictions.
+However, samples corresponding to certain feature values are often too sparse to learn an unbiased and confident calibrating network.
 As shown in Fig.
@@
 Most existing methods mainly focus on single field calibration, and it is impractical to directly merge multiple fields due to the ultra sparsity of samples satisfying multiple field value conditions.
-In this paper, we propose a confidence-aware multi-field calibration method named ConfCalib, which can perform fine-grained calibration on multiple feature fields meanwhile keeping awareness of the confidence degree of field data observation to adjust the calibration intensity adaptively.
-Compared with methods based on neural networks, ConfCalib, as a non-parametric method, performs more robustly in scenarios with severe data sparsity and has low costs of computation and deployment.
-Concretely, we regard user feedback on each field value as a binomial distribution and use Wilson intervals to compute confidence-aware deviation degrees of model predictions, where larger deviations indicate greater discrepancies between observed and predicted probabilities.
-We devise a mapping function to derive a smaller deviation and transform the predicted scores to achieve this new deviation.
-To further eliminate the influence of data sparsity, we develop a simple yet effective multi-field fusion method to synthesize the calibration results on different fields to enhance calibration performance and better accommodate advertising demands.
-Extensive offline experiments in different scenarios and the online A/B test on our advertising platform demonstrate the effectiveness of our approach.
+In this paper, we propose a confidence-aware multi-field calibration method named ConfCalib, which can perform fine-grained calibration on multiple feature fields and meanwhile keep robustness under data sparsity via adaptive adjustment of the confidence level.
+Concretely, we assume that user feedback follows a binomial distribution and use Wilson intervals to compute confidence-aware calibrated scores with a dynamic adjustment of calibration intensity.
+Moreover, we propose a simple yet effective multi-field fusion method to synthesize the calibration results on different fields, in order to further eliminate the influence of data sparsity.
+Compared with methods based on neural networks, our ConfCalib performs more robustly in scenarios with severe data sparsity.
+Extensive offline experiments in different datasets and online A/B test on our advertising platform demonstrate the effectiveness of our method.
 The main contributions of this paper are summarized as follows:
@@
   • Experiments on different offline datasets and the A/B testing on our advertising platform verify the superiority of our method.
-Figure 2. The overall framework of our ConfCalib method.
 2. Related Work
 Model calibration is a classic problem in machine learning, which is proposed to eliminate the discrepancy between the predicted probabilities of classification models and the true likelihood (Wald et al., 2021; Pleiss et al., 2017; Nixon et al., 2019).
-It has been widely explored due to the requirement of accurate probability prediction in various applications, such as ad auctions (He et al., 2014; McMahan et al., 2013), weather forecasting (Gneiting and Raftery, 2005; DeGroot and Fienberg, 1983), and personalized medicine (Jiang et al., 2012).
+It has been widely explored to promote accurate probability prediction in various applications, such as ad auctions (He et al., 2014; McMahan et al., 2013), weather forecasting (Gneiting and Raftery, 2005; DeGroot and Fienberg, 1983), and personalized medicine (Jiang et al., 2012).
 Calibration methods calibrate trained prediction models during training or inference phase.
@@
 Besides, label smoothing, which can be interpreted as a modified primary loss, also serves to model calibration (Müller et al., 2019; Pereyra et al., 2017; Liu et al., 2022).
-For model calibration in recommendation systems, post-hoc calibration methods are more prevalent.
-These methods transform the output of trained prediction models in the post-processing phase.
-According to the characteristics of different mapping functions, existing calibration methods can be divided into three categories: scaling-based, binning-based, and hybrid.
+This paper focus on post-hoc methods that calibrate model predictions by post-processing.
+According to the characteristics of different mapping functions, existing calibration methods can be divided into three categories: scaling-based, binning-based, and hybrid methods.
 Scaling-based methods (Guo et al., 2017; Mozafari et al., 2018; Kull et al., 2019; Platt et al., 1999; Kweon et al., 2022; Kull et al., 2017; Rahimi et al., 2020; Gupta et al., 2021) fit parametric mapping functions for the predicted scores.
@@
 To maintain the original order of model predictions, Isotonic Regression (Zadrozny and Elkan, 2002) adjusts the binning bounds by minimizing the residual between calibrated predictions and labels, hence ensuring monotonicity of the posterior probability in bins.
-Hybrid methods (Kumar et al., 2019; Zhang et al., 2020; Deng et al., 2021; Huang et al., 2022; Pan et al., 2020; Wei et al., 2022) combine scaling and binning to fit various mapping functions for different sample subsets, which allows a more precise calibration than single scaling or binning methods.
+Hybrid methods (Kumar et al., 2019; Zhang et al., 2020; Deng et al., 2021; Huang et al., 2022; Pan et al., 2020; Wei et al., 2022) combine scaling and binning methods, which are superior in calibration performance and prevalence in industry.
 Scaling-binning (Kumar et al., 2019) first fits a mapping function for the sorted samples and then employs histogram binning calibration on the mapped scores.
 Smoothed Isotonic Regression (SIR) (Deng et al., 2021) uses isotonic regression on bins and fits a monotonic piece-wise linear calibration function.
-Multiple Boosting Calibration Trees (MBCT) (Huang et al., 2022) optimizes the binning scheme through a tree-like division along feature fields and adopts linear scaling on each tree node individually.
 Recently, some works have focused on field-wise calibration, which aims to reduce the calibration error in a specific feature field.
@@
 AdaCalib (Wei et al., 2022) extends the single global mapping function of Neural Calibration to different ones for each value of a certain feature field and dynamically adjusts the number of bins by neural networks.
-These neural methods may suffer from data sparsity in specific application scenarios because networks are difficult to train.
+These neural methods may suffer from data sparsity in specific application scenarios where network training is hard to converge.
 Additionally, existing field-wise calibration methods simply consider one field and ignore potential relations among fields, meaning there is still a lack of research on multi-field calibration methods.
 3. Methodology
+Figure 2. The overall framework of our ConfCalib method.
 The overall framework of our multi-field confidence-aware calibration method is shown in Fig.
@@
 It first performs confidence-aware calibration on multiple target feature fields independently and then fuses the scores calibrated on different feature fields into a unified one.
-The details of our approach are introduced as follows.
+The details of our method are introduced as follows.
 3.1. Problem Definition
@@
 The above metrics will be detailed in Section 4.1.4.
+Figure 3. The curves of calibrated scores w.r.t. different predicted scores under different numbers of observed samples.
 3.2. Confidence-aware Calibration
-Figure 3. The curves of calibrated scores w.r.t. different predicted scores under different numbers of observed samples.
 Next, we introduce our confidence-aware calibration method.
-In practical scenarios, field-level calibration faces severe data sparsity issues due to the limited time window for data collection and the intrinsic sparsity of certain user feedback, such as conversions.
-Consequently, calibrations on fields with numerous feature values are neither accurate nor confident.
+In practical scenarios, field-level calibration faces severe data sparsity issues due to the limited time window for data collection and intrinsic sparsity of certain user feedback, such as conversions.
+Consequently, calibrations on feature fields with numerous values are neither accurate nor confident.
 The main problem caused by data sparsity is that the observed probability is not credible enough to reflect the true data distribution.
-In this case, complete trust in limited observations will likely result in overfitting.
-Therefore, it is necessary to preserve some trust in model predictions while referencing observations under data sparsity.
-Actually, many types of user feedback, like clicks and conversions, can be naturally modeled as binomial distributions.
-Given the observed positive sample ratio $p$ and the number of total samples $n$, we can then use a confidence interval formula to compute the confidence interval ($p_{-}$, $p_{+}$) of the real feedback probability under a certain confidence level.
-Conversely, given a certain predicted feedback probability, we can obtain the corresponding confidence level by considering the predicted value as the bound of the confidence interval.
+In this case, complete trust in limited observations will result in overfitting and harm generalization.
+In summary, it is necessary to preserve some trust in model predictions while referencing observations under data sparsity.
+Most user feedback, like clicks and conversions, can be naturally modeled using binomial distributions.
+Given the observed positive sample ratio $p$ and the number of total samples $n$, we can use a confidence interval formula to compute the confidence interval (CI) ($p_{-}$, $p_{+}$) of the real feedback probability under a certain confidence level.
+Conversely, given a certain predicted feedback probability, we can obtain the corresponding confidence level by regarding the predicted value as the bound of CI.
 In other words, there is a mapping between the predicted probability and the confidence level, which is the core of our confidence-aware calibration method.
-The most commonly used confidence interval formula is the normal interval.
+The most commonly used CI formula is the normal interval.
 However, due to the limitation of the central limit theorem, it is unreliable with a small sample size or an extreme probability close to 0 or 1, which is quite common for user feedback prediction in recommender systems.
@@
 where $z$ is the deviation score (e.g., $z=1.96$ for 95% confidence).
-In order to perform field-level calibration, our approach operates at the level of sample subsets, which can be obtained by splitting samples based on different values of the target feature field.
+In order to perform field-level calibration, our method operates at the level of sample subsets, which can be obtained by splitting samples based on different values of the target feature field.
 Besides, field-wise binning [^1] Binning is optional in our method since it introduces heavier data sparsity., i.e., further dividing each subset into several bins by order of their predicted scores, is optional.
 For each sample subset, we need the statistics, including the number of samples $n$, the ratio of positive samples $p$, and the average predicted score $\hat{p}$.
-By substituting these variables into Eq.
-(1), we can solve the value of the deviation score $z\in[0,+\infty)$ using numerical algorithms, such as the bisection method.
-A larger $z$ indicates greater discrepancies between model predictions and current observations.
-To calibrate the predictions, we propose to use a non-linear transformation function $g(z)$ to reduce the deviation.
-Intuitively, this function should satisfy the following requirements: (1) passing through the origin: $g(0)=0$; (2) monotonic: $\forall z_{1}\leq z_{2},g(z_{1})\leq g(z_{2})$; (3) bounded: $\exists C\in\mathbb{R}^{+},g(+\infty)\leq C$.
+Regard $\hat{p}$ as a certain bound [^2] Let $p_{+}=\hat{p}$ if $\hat{p}>p$, else $p_{-}=\hat{p}$.
+of CI and substitute these variables into Eq.
+(1), we can solve the deviation score $z\in[0,+\infty)$ using numerical algorithms, such as the bisection method.
+A larger $z$ indicates a greater bias between model predictions and current observations, but a higher confidence level as well.
+We can directly use $p$ as the calibrated score, i.e., change $z$ to 00, to entirely eliminate biases.
+However, this will minimize the confidence level and result in poor generalization under data sparsity.
+Therefore, we intend to make the calibrated score between original $p$ and predicted $\hat{p}$, thus corresponding to a relatively high confidence.
+To this end, we propose a non-linear transformation function $g(z)$ to reduce the deviation.
+Intuitively, $g(z)$ should satisfy the following requirements: (1) passing through the origin: $g(0)=0$; (2) monotonic: $\forall z_{1}\leq z_{2},g(z_{1})\leq g(z_{2})$; (3) bounded: $\exists C\in\mathbb{R}^{+},g(+\infty)\leq C$.
 Thus, we devise a variant of the sigmoid function as follows:
 (2) $\small g(z)=\lambda\cdot\left(\frac{2}{1+e^{-z/2}}-1\right),z\in[0,+\infty),$
-where $\lambda$ is a hyperparameter used to provide an upper bound and control the calibration intensity manually for large z-scores, i.e., deviations.
-A large calibration intensity means more belief in observations and less belief in original predictions.
-Obviously, we have $g(+\infty)=\lambda$ so that large deviations are calibrated to a relatively small one (e.g., the deviation corresponding to 95% confidence level if $\lambda=1.96$), and minor deviations are reduced almost linearly.
+where $\lambda$ is a hyperparameter used to provide an upper bound and control the calibration intensity manually.
+A larger calibration intensity can result in a smaller $z$, meaning more belief in observations and less belief in original predictions.
+Obviously, we have $g(+\infty)=\lambda$ so that a large deviation is calibrated to a small one (e.g., the deviation corresponding to 95% confidence level if $\lambda=1.96$), and minor deviations are reduced almost linearly.
 By replacing the original $z$ in Eq.
-(1) with the transformed value $g(z)$, we can obtain the calibrated average probability score $\hat{p}^{\prime}$.
-In order to ensure that $\hat{p}^{\prime}$ is located between the observed probability $p$ and the predicted probability $\hat{p}$, we use the positive branch if $\hat{p}$ is greater than $p$.
-Otherwise, we use the negative one.
+(1) with the transformed value $g(z)$, we can obtain a new CI, then one of its bounds is the desired confidence-aware calibrated score $\hat{p}^{\prime}$.
 In this way, given a test sample located in the same subset, we compute a field-wise calibrated score $p_{\text{calib}}$ by scaling its original predicted probability $p_{\text{pred}}$, i.e.,
@@
 Overall, our confidence-aware method adjusts model-predicted probabilities by resetting the confidence level, which considers the observation and prediction together to avoid violent value changes and overfitting limited observations.
-Consequently, it keeps robust when handling severe sample sparsity, such as circumstances with few observed samples and zero positive samples.
+Consequently, it keeps robust under severe data sparsity, especially circumstances with few observed samples and zero positive samples.
 3.3. Multi-field Joint Calibration
 Previous field-level calibration methods only calibrate and evaluate errors in a certain field, ignoring other fields.
-Actually, the field-level calibration error in different fields may vary greatly even for the same dataset, indicating a natural discrepancy of prediction biases in different fields.
+However, the field-level calibration error in different fields may vary greatly, which is ignored by single-field calibration.
 In addition, the biases in different fields, as reflections of the global bias from different perspectives, have a certain degree of commonality.
-Therefore, the data sparsity problem faced by a certain field is expected to be alleviated through calibrations on other fields.
-Therefore, simultaneously capturing biases of multiple fields can be significant to eliminate overall prediction bias and improve calibration performance.
+Therefore, data sparsity faced by a certain field is expected to be alleviated through calibrations on other fields.
 A naive way to extend calibration from a single field to multiple fields is to split sample subsets based on the combination of feature values of all the fields.
-Nevertheless, it is infeasible for the severer data sparsity since fewer samples satisfy multiple field value conditions.
-This means prior field-wise calibration methods for a single field cannot be directly extended to multi-field calibration.
-On the contrary, our method demonstrates better extensibility among fields.
-Note that the confidence-aware calibration of a certain field in Section 3.2 is actually a scaling operation for a specific sample and thus can be easily generalized to a multi-field joint calibration method.
+Nevertheless, it will cause severer data sparsity since fewer samples satisfy multiple field value conditions.
+Accordingly, prior single-field calibration methods cannot be directly extended to multi-field calibration.
+In contrast, our method demonstrates a better extensibility among fields.
+The confidence-aware calibration in a single field in Section 3.2 is actually a scaling operation for a specific sample, thus allowing a generalization to multi-field joint calibration.
 Assume that there are $K$ target fields to calibrate.
-Given a sample, we will get $K$ scaling multipliers $m_{1},m_{2},\dots,m_{K}$, i.e.
-${\hat{p}^{\prime}}/{\hat{p}}$ of each field.
-Then, a natural manner to adaptively fuse these multipliers is the weighted geometric mean,
+For each sample, we can obtain $K$ scaling multipliers $m_{1},m_{2},\dots,m_{K}$, where $m_{i}={\hat{p_{i}}^{\prime}}/{\hat{p_{i}}}$ for the $i$-th field.
+We adaptively fuse these multipliers via the weighted geometric mean, that is
 (4) $p_{\text{calib}}=m_{1}^{w_{1}}m_{2}^{w_{2}}\cdots m_{K}^{w_{K}}\cdot p_{\text{pred}},$
 where $w_{1},w_{2},\dots,w_{K}$ is weights satisfying $\sum_{i=1}^{K}{w_{i}}=1$.
-Considering the relatively small computation cost, the fusion weights can be precisely obtained by grid search.
-This method allows us to conduct joint calibration based on the marginal distribution of each field independently instead of dealing with the joint distribution, thus avoiding further sparsity of samples.
+The fusion weights can be precisely obtained by grid search.
+The proposed fusion method allows us to conduct joint calibration based on the marginal distribution of each field independently instead of dealing with the joint distribution, thus avoiding further sparsity of samples.
+Figure 4. Online deployment diagram of ConfCalib.
 3.4. Online Deployment
@@
 On the offline side, the prediction model’s training and ConfCalib’s fitting are paralleled and completely independent.
-At fixed intervals, the calibration module pulls the latest online samples from the log database and refits calibration curves.
+At fixed intervals, the calibration module pulls the latest online samples from the log database and refits calibration mappings.
 This interval is usually set relatively shorter (e.g., 30 minutes) than the prediction model update period to better capture the dynamic shift of online data.
 On the online side, prediction and calibration are conducted in sequence.
-Firstly, the prediction model makes predictions for candidate objects, then the calibration model calibrates the predicted probabilities and outputs the calibrated probabilities for subsequent use in the platform.
+First, the prediction model makes predictions for candidate objects, then the calibration model calibrates the predicted probabilities and outputs the calibrated probabilities for subsequent use.
 Recent methods designed for advertising platforms, such as NeuCalib and AdaCalib, involve auxiliary neural networks.
@@
 Therefore, it can be directly inserted into the pipeline without interfering with the training stage of prediction models, resulting in lower deployment costs.
-Figure 4. Online deployment diagram of ConfCalib.
 4. Experiments
@@
 For all the datasets and methods, the prediction model is a fully connected network with four layers.
-Hyperparameters of comparative methods, like the number of bins, are set according to the chosen value in the corresponding papers if provided.
-Otherwise, we tune the hyperparameters and select the optimal values.
-Considering the strong randomness of neural methods, we conduct five random experiments on each of them and report means and standard deviations of the results.
-For our method we set the number of bins as {50, 1, 10} and $\lambda$ in the z-score mapping function as {1.0, 0.2, 1.6}, corresponding to the following three datasets.
-Table 1. Detailed statistics of the three datasets.
+Hyperparameters of baselines are set according to the original papers or carefully tuned.
+Results of neural methods are the average results of five random experiments with means and standard deviations reported.
+For our ConfCalib, we set the number of bins as {50, 10} and $\lambda$ in Eq.
+(2) as {1.0, 1.6} for Avazu and our industrial dataset.
+Table 1. Data statistics of the public and industrial datasets.
 Dataset Task #Training #Validation #Test
 Avazu CTR 28.3M 4.04M 8.08M
-AliExpress CTCVR 62.2M 10.4M 31.2M
 Industrial CVR 27.5M 3.8M 7.5M
+4.1.2. Datasets
+We conduct experiments on our industrial dataset and a public dataset Avazu  [^3] https://kaggle.com/competitions/avazu-ctr-prediction.
+Data statistics are shown in Table 1.
+Avazu is a popular dataset on mobile ad CTR prediction with 40.4 million samples in 10 days ordered by the impression time.
+We split the dataset by day in a ratio of 7:1:2 for training, validation, and test.
+For our multi-field joint calibration method, the target fields are set to ”site_id” (3480) [^4] Represent the number of possible values of the corresponding field., ”app_id” (4864) and ”app_domain” (305).
+For other methods considering a single field, the target field is set to ”site_id”.
+The industrial dataset is collected from the click logs on our ad platform in 10 consecutive days, and the prediction target is conversion, with a split ratio of 7:1:2 by day.
+For our method, the target fields are ”advertiser_id” (529), ”app_size” (82), and ”app_id” (56).
+For other methods, the target field is ”advertiser_id”.
+The ratios of validation sets are relatively small while splitting, which is more in line with the short time window of online data collection.
+4.1.3. Baselines
+We choose several representative methods of the three categories as baselines for comparison.
+The scaling-based methods include naive scaling (multiply all predicted scores by a value $k$ to make their mean equal to the positive sample ratio), Platt Scaling (PlattScaling) (Platt et al., 1999), Gaussian Calibration (GaussCalib) (Kweon et al., 2022) and Gamma Calibration (GammaCalib) (Kweon et al., 2022).
+The binning-based methods include Histogram Binning (HistoBin) (Zadrozny and Elkan, 2001) and Isotonic Regression (IsoReg) (Zadrozny and Elkan, 2002).
+The hybrid methods include Smoothed Isotonic Regression (SIR) (Deng et al., 2021), Neural Calibration (NeuCalib) (Pan et al., 2020) and AdaCalib (Wei et al., 2022), with the latter two methods based on neural networks.
+The number of bins for NeuCalib and AdaCalib is set according to the initial settings when proposed.
+4.1.4. Evaluation Metrics
+Referring to prior methods, we continue to use field-level relative calibration error (Field-RCE or F-RCE) (Pan et al., 2020) to evaluate the calibration error on a certain field, which formula is
+[…]
+where $\mathcal{D}$ is the dataset, $\mathcal{Z}$ is the possible value set of the target field and $\mathcal{D}_{z}$ is the subset of samples with the target field value $z$ satisfying $\sum_{z\in\mathcal{Z}}{|\mathcal{D}_{z}|}=|\mathcal{D}|$.
+Intuitively, a lower Field-RCE indicates lower biases along the target field.
+Furthermore, we use the mean Field-RCE on multiple fields (Multi-Field-RCE or MF-RCE) to measure the multi-field calibration performance.
+Meanwhile, we adopt expected calibration error (ECE) (Naeini et al., 2015) and multi-view calibration error (MVCE) (Huang et al., 2022) to evaluate the global calibration error.
+ECE arranges all predicted values in order and divides them equally into several subsets, i.e., a binning operation.
+Then it is calculated as
+(6) $\small\text{ECE}=\frac{1}{|\mathcal{D}|}\sum_{t=1}^{T}{{\left|\sum\nolimits_{(x_{i},y_{i})\in\mathcal{D}_{t}}{(y_{i}-\hat{p}_{i})}\right|}},$
+where $T$ is the number of bins and $\mathcal{D}_{t}$ is the subset of samples in the $t$-th bin.
+MVCE iteratively shuffles the dataset and calculates ECE until the accumulative mean ECE converges.
+The formula is
+(7) $\small\text{MVCE}=\frac{1}{RT}\sum_{r=1}^{R}\sum_{t=1}^{T}{\frac{\left|\sum_{(x_{i},y_{i})\in\mathcal{D}_{r,t}}{(y_{i}-\hat{p}_{i})}\right|}{|\mathcal{D}_{r,t}|}},$
+where $R$ is the number of shuffle operations, $T$ is the number of bins, and $\mathcal{D}_{r,t}$ is the subset of samples in the $t$-th bin after the $r$-th shuffle.
+MVCE allows us to inspect the calibration error from a more global perspective beyond the level of feature fields, thereby highlighting the effectiveness of multi-field calibration.
+Additionally, we also report AUC and LogLoss to evaluate the ranking performance after calibration since our method and primary baselines are not globally isotonic.
 Table 2. The performance of different methods in terms of prediction accuracy and miscalibration degrees. $\uparrow$” means higher is better, $\downarrow$” means lower is better.
-Dataset Type Method AUCnormal-$\uparrow$ LogLossnormal-$\downarrow$ Field-RCEnormal-$\downarrow$ Multi-field-RCEnormal-$\downarrow$ ECEnormal-$\downarrow$ MVCEnormal-$\downarrow$
+[…]
 Avazu (CTR) No Calib. N/A 0.7545 0.3718 0.2295 0.1837 0.0123 0.0117
@@
 ConfCalib 0.7575 0.3693 0.1435 0.0967 0.0037 0.0015
-AliExpress (CTCVR) No Calib. N/A 0.8773 0.0046 0.3274 0.3273 2.085e-4 2.085e-4
-Binning HistoBin 0.8752 0.0047 0.1417 0.1411 8.688e-5 8.742e-5
-IsoReg 0.8770 0.0046 0.1397 0.1394 9.811e-5 8.568e-5
-Scaling Naive 0.8773 0.0046 0.1431 0.1429 8.379e-5 8.450e-5
-PlattScaling 0.8773 0.0045 0.1406 0.1403 8.597e-5 8.658e-5
-GaussCalib 0.8773 0.0045 0.1433 0.1430 8.784e-5 8.836e-5
-GammaCalib 0.8773 0.0045 0.1404 0.1401 8.538e-5 8.600e-5
-Hybrid SIR 0.8773 0.0046 0.3299 0.3298 2.102e-4 2.102e-4
-NeuCalib 0.8773$\pm$0.0003 0.0046$\pm$5e-7 0.1499$\pm$0.0480 0.1497$\pm$0.0480 1.068e-4$\pm$3.236e-5 8.617e-5$\pm$3.300e-5
-AdaCalib 0.8759$\pm$0.0004 0.0046$\pm$5e-8 0.1459$\pm$0.0383 0.1458$\pm$0.0375 1.045e-4$\pm$3.131e-5 8.003e-5$\pm$3.316e-5
-ConfCalib 0.8778 0.0045 0.1360 0.1360 8.352e-5 8.425e-5
 Industrial (CVR) No Calib. N/A 0.8226 0.0868 0.3657 0.3015 0.0022 0.0008
@@
 ConfCalib 0.8303 0.0860 0.2728 0.2243 0.0007 0.0007
-4.1.2. Datasets
-We conduct experiments on our industrial dataset, as well as two public datasets, including Avazu  [^2] https://kaggle.com/competitions/avazu-ctr-prediction and AliExpress (Li et al., 2020).
-The data statistics are shown in Table 1.
-Avazu is a popular dataset on mobile ad click-through rate (CTR) prediction with 40.4 million samples in 10 days ordered by the impression time.
-We split the dataset by day in a ratio of 7:1:2 for training, validation, and test.
-For our multi-field joint calibration method, the target fields are set to ”site_id” (3480) [^3] Represent the number of possible values of the corresponding field., ”app_id” (4864) and ”app_domain” (305).
-For other methods considering a single field, the target field is set to ”site_id”.
-AliExpress is a dataset for multi-task learning of click-through rate and conversion rate prediction in e-commerce.
-Here, we choose the data from four countries except Russia and only use the conversion label to build a single click-through conversion rate (CTCVR) prediction task.
-The training, validation, and test set are split in a ratio of 6:1:3, totaling 103.8 million samples.
-The timestamp of the test set is strictly later than the other two sets.
-For our method, the target fields are ”country” (4) and an anonymous field, ”feature_1” (10).
-For other methods, the target field is ”feature_1”.
-For the industrial data, we collect 38.8M clicked samples in 10 consecutive days from the log of our ad platform to build a post-click conversion rate (CVR) prediction task with a split ratio of 7:1:2 by day.
-For our method, the target fields for calibration are ”advertiser_id” (529), ”app_size” (82), and ”app_id” (56).
-For other methods, the target field is ”advertiser_id”.
-The ratios of validation sets are relatively small while splitting, which is more in line with the short time window for data collection during online running.
-4.1.3. Comparative Methods
-We choose several representative methods of the three categories of methods for comparison.
-The scaling-based methods include naive scaling (multiply all predicted scores by a value $k$ to make their mean equal to the positive sample ratio), Platt Scaling (PlattScaling) (Platt et al., 1999), Gaussian Calibration (GaussCalib) (Kweon et al., 2022) and Gamma Calibration (GammaCalib) (Kweon et al., 2022).
-The binning-based methods include Histogram Binning (HistoBin) (Zadrozny and Elkan, 2001) and Isotonic Regression (IsoReg) (Zadrozny and Elkan, 2002).
-The hybrid methods include Smoothed Isotonic Regression (SIR) (Deng et al., 2021), Neural Calibration (NeuCalib) (Pan et al., 2020) and AdaCalib (Wei et al., 2022), with the latter two methods based on neural networks.
-The number of bins for NeuCalib and AdaCalib is set according to the initial settings when proposed.
-4.1.4. Evaluation Metrics
-Referring to prior methods, we continue to use field-level relative calibration error (Field-RCE) (Pan et al., 2020) to evaluate the calibration error on a certain field, which formula is
-[…]
-where $\mathcal{D}$ is the dataset, $\mathcal{Z}$ is the possible value set of the target field and $\mathcal{D}_{z}$ is the subset of samples with the target field value $z$ satisfying $\sum_{z\in\mathcal{Z}}{|\mathcal{D}_{z}|}=|\mathcal{D}|$.
-Intuitively, a lower Field-RCE indicates lower biases along the target field.
-Furthermore, we use the mean Field-RCE on multiple fields (Multi-field-RCE) to measure the multi-field calibration performance.
-Meanwhile, we adopt expected calibration error (ECE) (Naeini et al., 2015) and multi-view calibration error (MVCE) (Huang et al., 2022) to evaluate the global calibration error.
-ECE arranges all predicted values in order and divides them equally into several subsets, i.e., a binning operation.
-Then it is calculated as
-(6) $\small\text{ECE}=\frac{1}{|\mathcal{D}|}\sum_{t=1}^{T}{{\left|\sum\nolimits_{(x_{i},y_{i})\in\mathcal{D}_{t}}{(y_{i}-\hat{p}_{i})}\right|}},$
-where $T$ is the number of bins and $\mathcal{D}_{t}$ is the subset of samples in the $t$-th bin.
-MVCE iteratively shuffles the dataset and calculates ECE until the accumulative mean ECE converges.
-The formula is
-(7) $\small\text{MVCE}=\frac{1}{RT}\sum_{r=1}^{R}\sum_{t=1}^{T}{\frac{\left|\sum_{(x_{i},y_{i})\in\mathcal{D}_{r,t}}{(y_{i}-\hat{p}_{i})}\right|}{|\mathcal{D}_{r,t}|}},$
-where $R$ is the number of shuffle operations, $T$ is the number of bins, and $\mathcal{D}_{r,t}$ is the subset of samples in the $t$-th bin after the $r$-th shuffle.
-MVCE allows us to inspect the calibration error from a more global perspective beyond the level of feature fields, thereby highlighting the effectiveness of multi-field calibration.
-The above calibration error metrics focus on different aspects, thus allowing us to examine the performance of each method more comprehensively.
-Additionally, we also report AUC and LogLoss to evaluate the ranking performance after calibration since our method and main comparative methods are not globally isotonic.
+Figure 5. Results of the ablation study on several variants of ConfCalib.
 4.2. Results
-Figure 5. Results of ablation studies on several variants of ConfCalib.
 4.2.1. Main Results
 The main results on each dataset are shown in Table 2.
-Our method shows better calibration performance compared with the competing methods.
-It achieves lower calibration errors on both field-wise and global metrics.
-Besides, our method also achieves higher AUC in all the tasks, indicating improvement in ranking ability in contrast to the uncalibrated model and other methods that may change the original order.
-By contrast, the performance of scaling and binning-based comparative methods is unsatisfactory.
-These methods reduce global errors to some extent through global calibration, but the performance is not outstanding.
-Meanwhile, they are basically powerless against field-wise calibration errors.
-Field-level calibration methods with neural networks such as NeuCalib and AdaCalib achieve relatively low calibration errors on the single target field among comparative methods.
+Our method shows better calibration performance compared with baselines.
+It achieves lower calibration errors on both field-wise and global metrics compared to the uncalibrated model and almost all the baselines.
+Besides, our method also achieves higher AUC, indicating an improvement in ranking ability.
+Scaling and binning-based baselines show poor performances.
+They reduce global errors to some extent through global calibration while powerless against field-wise calibration errors.
+Field-level calibration methods such as NeuCalib and AdaCalib achieve low calibration errors on the single target field among baselines.
 However, they are still inferior to our method, especially on multi-field metrics like Multi-Field-RCE and global metrics like ECE, demonstrating the effectiveness of our multi-field joint calibration.
-Moreover, the results of the two neural methods show large variance, which means significantly unstable learning processes on small calibration datasets.
-On the contrary, our method only relies on posterior statistical information and maintains a stabler calibration effect.
-It is noteworthy that results of the neural method AdaCalib on industrial data and AliExpress indicate that it may be unable to learn effectively in some cases, resulting in AUC even lower than the uncalibrated one.
-This phenomenon is likely caused by severe data sparsity.
-After sequential divisions of the feature value split and the binning operation, the number of total and positive samples within each subset is very limited.
-Thus, it is difficult for AdaCalib to learn the network parameters of corresponding parts of the piece-wise calibration curves.
-Our method works well in the same situation, benefiting from the awareness of confidence and the multi-field fusion approach designed to deal with sparse data.
+The two neural methods show large variance, which means unstable learning on sparse data.
+On the contrary, our method only relies on posterior statistical information and maintains a stable calibration effect.
+Moreover, AdaCalib obtains a lower AUC than the uncalibrated model on our industrial dataset, indicating that complex neural calibration methods may be counterproductive under data sparsity.
 4.2.2. Ablation Study
@@
 5.
-Concretely, we respectively remove the confidence-aware calibration with Wilson interval, the nonlinear z-score mapping function (replaced by a single linear mapping), and the multi-field joint calibration from the original ConfCalib.
-The results show that all variants have varying degrees of increase in calibration errors and decrease in AUC, which can demonstrate the effectiveness of each key point.
-The confidence-aware calibration partly preserves confidence in original predictions while calibrating based on observed posterior probabilities, which can avoid overfitting on limited validation samples in case the bias direction in the test dataset is opposite.
-In other words, it can effectively mitigate the impact of real-time data shifts during online serving.
-Although we cannot simulate frequent data shifts using a fixed test dataset, there is still an evident decline in AUC and an increase in Field-RCE after removing the confidence-aware calibration.
-The elaborate z-score mapping function dynamically adjusts the calibration intensity according to the deviations between predicted and observed probabilities.
-For samples with large deviations, it introduces an upper bound to force the calibrated result back to a high enough confidence level.
-After replacing the mapping function with a single linear function $g(z)=\lambda z$, AUC and all calibration error metrics significantly deteriorate.
-The main reason is that the unbounded linear function chooses insufficient calibration strength for samples with large deviations.
-The multi-field joint calibration simultaneously considers the biases across multiple fields and conducts joint calibration with optimized weights of each field.
-Compared to calibration on a single field, the joint calibration on multiple fields significantly improves both field-wise and global calibration errors, as well as AUC.
+Concretely, we respectively remove the confidence-aware calibration (i.e., turn into a field-wise naive calibration), the nonlinear z-score mapping function (replaced by a single linear mapping), and the multi-field joint calibration from the original ConfCalib.
+The removal of our confidence-aware calibration results in increase in field-wise calibration errors and decrease in AUC.
+After replacing the mapping function with a single linear function $g(z)=\lambda z$, AUC and all calibration error metrics deteriorate.
+This validates that the bounded mapping can provide proper calibration intensity compared with simple scaling.
+Finally, removing the joint calibration on multiple fields significantly increases both field-wise and global calibration errors and causes a large decline in AUC.
 Figure 6. Results of robustness analysis of different data sparsity levels.
-Table 3. Results of recalibration on comparative methods.
-Method AUCnormal-$\uparrow$ Field-RCEnormal-$\downarrow$ Multi-field-RCEnormal-$\downarrow$ ECEnormal-$\downarrow$ MVCEnormal-$\downarrow$
-GaussCalib
-0.7545
-0.1969
-0.1470
-0.0058
-0.0041
-+ConfCalib 0.7577 0.1417 0.1105 0.0053
-0.0050
-GammaCalib
-0.7545
-0.1984
-0.1480
-0.0058
-0.0041
-+ConfCalib 0.7577 0.1430 0.1113 0.0053
-0.0050
-NeuCalib
-0.7554
-0.1491
-0.1091
-0.0110
-0.0042
+Table 3. Results of recalibration on learning-based baselines.
+[…]
+GaussCalib 0.7545 0.1969 0.1470 0.0058 0.0041
++ConfCalib 0.7577 0.1417 0.1105 0.0053 0.0050
+GammaCalib 0.7545 0.1984 0.1480 0.0058 0.0041
++ConfCalib 0.7577 0.1430 0.1113 0.0053 0.0050
+NeuCalib 0.7554 0.1491 0.1091 0.0110 0.0042
 +ConfCalib 0.7558 0.1446 0.0994 0.0099 0.0030
-AdaCalib
-0.7569
-0.1650
-0.1308
-0.0099
-0.0065
-+ConfCalib
-0.7569
-0.1535 0.1069 0.0072 0.0024
-4.3. Recalibration on Comparative Methods
-Part of the comparative methods are learning-based, such as GaussCalib, GammaCalib, NeuCalib, and AdaCalib.
-It can be seen from the results in Table 2 that the performances of these methods are not satisfying for remaining high calibration errors or reducing AUC.
-Since our method does not require extra model training and has a low deployment cost, it can recalibrate the calibrated output of other methods for further improvement.
+AdaCalib 0.7569 0.1650 0.1308 0.0099 0.0065
++ConfCalib 0.7569 0.1535 0.1069 0.0072 0.0024
+4.3. Recalibration on Learning-Based Baselines
+Part of the baselines are learning-based, such as GaussCalib, GammaCalib, NeuCalib, and AdaCalib.
+Since our ConfCalib does not require extra model training and has a low deployment cost, it can recalibrate the calibrated output of other methods for further improvement.
 We experiment on Avazu CTR task.
-In detail, we first independently fit our method and the comparative methods to be recalibrated on the validation set.
-Then, the comparative methods make calibrated predictions for the test set.
+First, we fit our method and baselines to be recalibrated separately on the validation set.
+Then, baselines make calibrated predictions for the test set.
 Finally, our method conducts recalibration on the calibrated outputs.
-Table 3 shows that recalibration using our ConfCalib achieves further improvement compared with the original results of other calibration methods.
-Profiting from the multi-field joint calibration, the recalibration significantly reduces Multi-field-RCE and the global calibration error MVCE.
-Besides, there is a large increase in AUC for scaling-based methods, including GaussCalib and GammaCalib, which strictly keep the order of the original predictions.
+Table 3 shows that recalibration achieves further improvement compared with the original results of other methods.
+Profiting from multi-field joint calibration, recalibration using ConfCalib significantly reduces Multi-Field-RCE and the global calibration error MVCE.
+Besides, recalibration on methods that strictly keep the order of original predictions, such as GaussCalib and GammaCalib, can provide an improvement in AUC as well.
 4.4. Robustness Analysis against Data Sparsity
-As described in Section 3.2, one main advantage of ConfCalib is its robustness against data sparsity.
-Due to the limited time window for online data collection, the number of samples available to fit calibration models can be very few.
-Under such circumstances, methods based on neural networks can lead to inferior effects, although they may behave well if trained on sufficient samples.
-Observing the results on Avazu in Table 2, we can find that the neural methods NeuCalib and AdaCalib both achieve lower calibration errors compared with the original predictions.
-Here, we reduce the number of validation samples in Avazu with various sample rates to simulate scenarios with varying degrees of data sparsity.
+One primary advantage of ConfCalib is its robustness against data sparsity.
+Due to the limited time window for online data collection, samples available to fit calibration models can be very few.
+Under such circumstances, methods based on neural networks may be ineffective or even counterproductive.
+Here, we reduce the validation dataset used for calibration in Avazu with various sample rates to simulate scenarios with varying degrees of data sparsity.
 The calibration performances w.r.t different sparsity degrees are shown in Fig.
 6.
-As the number of samples for fitting the calibration model decreases, the performances of NeuCalib, AdaCalib, and ConfCalib all deteriorate in general.
-Fluctuations in the results of the two neural methods are intense, indicating that they are susceptible to data sparsity.
-However, our ConfCalib is little affected and exhibits strong robustness against data sparsity.
+As the number of samples for fitting the calibration model decreases, the performances of NeuCalib, AdaCalib get worse evidently, while our ConfCalib is little affected and exhibits strong robustness against data sparsity.
+Figure 7. Fusion weights in multi-field joint calibration.
 4.5. Analysis of Multi-field Fusion Weights
@@
 Nevertheless, when Field-RCE on ”ad_type” is minimized, the weight of the other field ”consume_frequency” is the highest instead.
-This phenomenon indicates that calibration on a single field may not be sufficient, while calibration on other fields can provide effective assistance.
-Figure 7. Fusion weights in multi-field joint calibration.
+This indicates that calibration on a single field is insufficient, and multi-field calibration can provide effective assistance.
 4.6. Hyperparameter Analysis
-Our method has a single hyperparameter $\lambda$ located in the z-score mapping function to control the calibration strength manually.
+Our method has a single hyperparameter $\lambda$ in the z-score mapping function Eq.
+(2) to control the calibration strength manually.
 Its main effect is to select a relatively higher confidence level when observed samples are adequate, thus making the calibrated prediction closer to the observation.
 Here, we discuss the influence of different settings of $\lambda$.
-In detail, we set $\lambda\in\{0,0.2,0.4\dots,3.0\}$ and perform experiments on the industrial dataset.
-The results are shown in Fig.
+We set $\lambda=0,0.2,0.4\dots,3.0$ and experiment on the industrial dataset.
+Results are shown in Fig.
 8.
-Overall, the fluctuation of field-wise calibration errors like Field-RCE and Multi-field-RCE is within a small range as $\lambda$ changes.
-For global calibration error metrics, the change of ECE is relatively violent when $\lambda$ is large, while MVCE remains almost unchanged throughout.
+Overall, the variation of field-wise calibration errors like Field-RCE and Multi-Field-RCE is within a small range as $\lambda$ changes.
+For global calibration error metrics, the change of ECE is relatively violent when $\lambda$ is significant, while MVCE remains almost the same.
 When $\lambda=0$, the awareness of confidence is removed, which means we disregard the model prediction and believe entirely in the observation.
 We achieve a local optimum at this point.
-As $\lambda$ increases, the calibration intensity gradually decreases, and the global optimal $\lambda$ is around 1.6, considering all the calibration errors.
+As $\lambda$ increases, the calibration intensity gradually decreases, and the global optimal $\lambda$ is around 1.6.
 Figure 8. Calibration errors w.r.t hyperparameter $\lambda$.
 4.7. Online A/B Test
-We conduct one week’s online A/B test separately on two business scenarios to evaluate the effectiveness of ConfCalib in practical applications.
-In each scenario, we randomly split users into two groups, each with tens of millions of users everyday.
-One group receives recommendations by ConfCalib, while the other one gets recommendations by a well-crafted baseline.
-On an industrial advertising platform, ConfCalib achieves a relative improvement of 2.42% in CVR compared with the baseline.
-On the Top-grossing of a popular Android app store, ConfCalib obtains relative improvement of 32.6% and 49.1% in CTR and the revenue over the baseline.
-The data distribution shifts frequently on the Top Grossing scenario, and ConfCalib can better adapt this variability to achieve a significant improvement.
+We conduct one week’s online A/B test separately on two industrial scenarios to evaluate ConfCalib’s effectiveness in practical applications.
+In each scenario, we randomly split users into two groups, each with tens of millions of users daily.
+One group receives recommendations from ConfCalib, while the other gets recommendations from a well-crafted baseline.
+On Huawei’s online advertising platform, ConfCalib achieves a relative improvement of 2.42% in CVR compared with the baseline.
+On Top-grossing of Huawei AppGallery, ConfCalib obtains relative improvement of 32.6% and 49.1% in CTR and revenue.
+Top-grossing is a small-scale scenario where data distribution shifts frequently, and ConfCalib can better adapt this variability to achieve a significant improvement.
 5. Conclusion
-In this paper, we propose a confidence-aware multi-field model calibration method, ConfCalib, to address the common phenomenon of data sparsity in model calibration on advertising platforms.
-ConfCalib can dynamically adjust the calibration intensity based on the number of observed samples and posterior probabilities while considering the biases along multiple feature fields for joint calibration to capture various bias patterns and further alleviate the impact of data sparsity.
-We conduct sufficient comparative experiments on public and industrial datasets of different user action prediction tasks, and the results demonstrate that ConfCalib outperforms prior methods and exhibits strong robustness against data sparsity.
-ConfCalib is currently running on an industrial advertising platform and a popular Android app store, and has achieved continuous online revenue improvement.
+In this paper, we propose a confidence-aware multi-field model calibration method, ConfCalib, to address the common phenomenon of data sparsity in model calibration for advertising recommendation.
+ConfCalib dynamically adjusts calibration intensity based on the confidence level obtained from the observed data distribution.
+Moreover, it considers biases along multiple feature fields and introduces joint calibration to capture various bias patterns and further alleviate the impact of data sparsity.
+We conduct sufficient offline and online experiments.
+Results demonstrate that ConfCalib outperforms prior methods and exhibits strong robustness against data sparsity.
+ConfCalib is currently running on Huawei’s advertising platform and AppGallery and has achieved continuous online revenue improvement.